# Refine Neutrino Events Reconstruction with BEiT-3

### v2

@@ III.3 Multilayer Perceptron: Regression and Classification
For a continuous space, regression and classification are commonly used methods to predict the output of a neural network. We use different Multilayer Perceptron (MLP) as the final layer to build these two models:
@@
  • Regression: The model predicts the output itself. The output is produced by a MLP that results with a tensor with shape $[B,3]$, where $B$ is mini-batch size and 3 represents $x_{pred}$, $y_{pred}$, $z_{pred}$, which belong to $[-1,1]$.
@@
We trained three models (The hyper-parameter is summarized in table 3. For detailed training configuration and loss curve, please see section IV.2.):
@@
  • Model S-RegB is a regression model using similar structure, replacing one encoder with two encoders with and without Relative Position Bias, and residual connections are used bewteen these two encoders.
@@
III.4 Loss
We trained the regression model using Mean Squared Error (MSE) Loss, and the classification model using Cross Entropy (CE) Loss. The competition Metric is used as metric in the validation process.
III.4.1 Mean Squared Error Loss
As stated in section II.2, the shape of IceCube is close to a cylinder, and the first Bessel function is the solution for partial differential equations in a cylinder, thus the Von Mises-Fisher Loss works really well. But in ISeeCube, a more primitive loss function such as Mean Squared Error (MSE) Loss can work just well.
Mean Squared Error (MSE) Loss is a common loss function for regression models. It calculates mean squared error between each element in the prediction $x$ and target $y$:
$\textrm{MSE}\left(x,y\right)=\frac{1}{n}\sum_{i}\left(x_{i}-y_{i}\right)^{2}$ (2)
where $x_{i}$ and $y_{i}$ can be in any shape. In our case, $x_{i}$ and $y_{i}$ are both one-dimensional array with length $3$ (representing $x_{\text{reco}}$, $y_{\text{reco}}$, $z_{\text{reco}}$ and $x_{\text{true}}$, $y_{\text{true}}$, $z_{\text{true}}$ respectively).
Since MSE Loss is a rather primitive loss function, it could be used for more tasks such as energy reconstruction. And to prevent exploding gradients, it can be replaced by Smooth L1 Loss Girshick (2015).
III.4.2 Cross Entropy Loss
And for classification models, following the Kaggle solutions stated in II.2, we use another rather primitive loss function Cross Entropy (CE) Loss:
$\textrm{CE}\left(x,y\right)=\frac{1}{n}\sum_{i}x_{i}\log{y_{i}}$ (3)
@@ III.4.2 Cross Entropy Loss
Similarly, since CE Loss is a rather primitive loss function, it could be used for more tasks such as cascade/track classification.
III.4.3 Competition Metric
The derivation of Competition Metric Eller (2023) (Mean Angular Error) is as follows. In spherical coordinate system, given two vectors true direction $\vec{r}_{\textrm{true}}$ and predicted direction $\vec{r}_{\textrm{reco}}$ with radius equals $1$ (after normalization), we have:
$\vec{r}_{\text{true}}$ $=\hat{r}+\hat{\theta}\vartheta_{\text{true}}+\hat{\phi}\varphi_{\text{true}}$ (4)
$\vec{r}_{\text{reco}}$ $=\hat{r}+\hat{\theta}\vartheta_{\text{reco}}+\hat{\phi}\varphi_{\text{reco}}$
In 3D Cartesian coordinate system, after normalization (with $r_{\textrm{reco}}=1$):
$x_{\text{reco}}$ $=\sin{\vartheta_{\text{reco}}}\cos{\varphi_{\text{reco}}}$ (5)
$y_{\text{reco}}$ $=\sin{{\vartheta_{\text{reco}}}}\sin{{\varphi_{\text{reco}}}}$
$z_{\text{reco}}$ $=\cos{\vartheta_{\text{reco}}}$
these two vectors are transformed into:
$\vec{r}_{\text{true}}$ $=\hat{i}\sin{\vartheta_{\text{true}}}\cos{\varphi_{\text{true}}}+\hat{j}\sin{{\vartheta_{\text{ture}}}}\sin{{\varphi_{\text{true}}}}+\hat{k}\cos{\vartheta_{\text{true}}}$ (6)
$\vec{r}_{\text{reco}}$ $=\hat{i}\sin{\vartheta_{\text{reco}}}\cos{\varphi_{\text{reco}}}+\hat{j}\sin{{\vartheta_{\text{reco}}}}\sin{{\varphi_{\text{reco}}}}+\hat{k}\cos{\vartheta_{\text{reco}}}$
Thus the angular error is the arccosine of the inner product of these two vectors:
$\Psi=\arccos(\sin{\vartheta_{\text{true}}}\sin{\vartheta_{\text{reco}}}(\cos{{\varphi_{\text{ture}}}}\cos{{\varphi_{\text{reco}}}}+\sin{{\varphi_{\text{ture}}}}\sin{{\varphi_{\text{reco}}}})+\cos{\vartheta_{\text{true}}}\cos{\vartheta_{\text{reco}}})$ (7)
And Mean Angular Error on the entire test dataset is the average of $\Psi$ for every event in the entire dataset. Random guessing will get a result of $\frac{\pi}{2}$, when two vectors are orthogonal.
IV Experiments
In section IV.1 we summarize some of the most important aspects of the dataset and the normalization of neutrino events. In section IV.2 we show the learning rate schedule, loss curve and performance of our models, followed by section IV.3 we further explore the performance of our models by comparing azimuthal and zenithal distribution. Also in this section we propose a new metric: Overlap Ratio.
IV.1 Dataset
Each event in the dataset is composed of a series of pulses. Most neutrino events have less or equal to $200$ pulses, and in most Kaggle solutions the dimension of the model is less than $200$, meaning that if the number of pulses is bigger than the dimension, some of the pulses will be discarded.
@@ IV.1 Dataset
Figure 3: This is an example event where auxiliary False and True are shown separately. It is intuitive that False will get better result in terms of shape thus direction reconstruction. This figure is originally used in Eller (2023).
The biggest challenge of the dataset is its bigness (more than $138$ million events distributed in $660$ batches with the size of over 100 GBs), so we use the same chunk-based training method used in the $2^{\textrm{nd}}$ place solution, meaning that every epoch is trained on $654$ batches divided into $8$ sub-epochs, see section IV.2. The advantage of this chunk-based method is that now we can train the model on the entire dataset with relatively low memory cost. However it may have a disadvantage that a random sampling can only sampling across the sub-set of the entire dataset, which may leads to bias.
IV.2 Training and Performance
@@
Table 3: Hyper-parameters and score of ISeeCube. “S-RegA”, “S-RegB” and “S-ClsB” refers to different “Small” models (see section III.3). The score for competition metric is the smaller the better.
@@
Model S-RegA follows the learning rate schedule as table 4. Model S-RegB and S-ClsB follows the learning rate schedule as table 5. The reason for epoch 0 in table 4 is that we first train our model for 1 epoch to test the performance and we think load this pre-trained model for further training won’t lead to overfitting, considering that (a) the score is relatively high and (b) there are following 7 epochs with different dataset (see chunk-based method mentioned in section IV.1). In order to save time and resources when training, for epoch 0$\sim$48, we use $N=196$ when loading the data (the pulses in an event is 196 after normalization), and for epoch 49$\sim$56, we use $N=256$. In table 4 and table 5, “COMB” loss means combining two losses by adding them up:
$COMB$ $=\Psi+0.05\cdot MSE$ (8)
@@
where $\Psi$ is competition metric in the Kaggle competition. “div” and “div_final” are hyper-parameter used in Leaner.fit_one_cycle with 1cycle police Smith and Topin (2019) in fastai.
@@
The loss curve of S-RegA, S-RegB and S-ClsB are shown in figure 4. As we can see, the training loss curve is flucuating while the validation loss curve is constantly reducing. Also, by replacing MSE loss with COMB loss, the loss curve drops in a single epoch.
@@
As shown in table 2 and table 3, by adding merely 2.4M parameters, our best model S-RegA outperforms $2^{\textrm{nd}}$ place solution’s S d32 model and approaches its B d32 model. Also, compared with CV512, our model uses CV256, which means that our model requires half the GPU memory when inferencing.
Figure 5: By using voting rule as the ensemble method, our model reaches a score of 0.9645.
IV.3 Azimuthal and Zenithal Distribution
@@
$\rho=\frac{\text{overlap area of the prediction and the true value}}{\text{area of the true value}}$ (9)
As shown in figure 6, the overlap ratio of both regression model and classification model are over 90%.
@@
On the contrary, the zenithal distribution is irregular and has lower overlap ratio. The predictions made by regression models tend to be drifted towards $\theta=0$, while the prediction made by classification tends to be drifted towards $\theta=\frac{\pi}{3}$ and $\theta=\frac{\pi}{2}$. Regression models tend to have a higher overlap ratio than classification models.
@@
As we can see from table 3 and figure 6, this ratio uniformly reflects the performance of the model. By fitting the score of the competition metric and the overlap ratio with four data points (S-RegA, S-RegB, S-ClsB and ensemble), we can get figure 7. The curve for zenith ($\theta$) is more linear than the curve for azimuth ($\varphi$), this is most likely due to the fact that azimuth distribution is more uniform than zenith distribution, as shown in figure 6.
@@
We use some of the most common loss in machine learning to investigate the fundamental models for Neutrino Events Reconstruction. The bias of two major models (Regression and Classification) is like two major interpretations of black-body radiation (Wien’s law and Rayleigh–Jeans law), meaning that each prediction has its own bias and we are still looking for a unified model. By using ensemble, we can lower the competition metric score, but the distribution of zenith is largely close to regression. Thus regression is not the unified model.
@@
Figure 7: Fitting competiton metric score with overlap ratio. We can see that overlap ratio uniformly goes up with better competition metric score.
V Discussion and Conclusion
@@
Neutrino Events Reconstruction has always been crucial for neutrino observatories such as IceCube and KM3NeT/ORCA, which traditionally uses CNN or GNN. In the Kaggle competition, many solutions use Transformer. And it is shown that a pure Transformer model performs really well for Neutrino Events Reconstruction.
@@
Compared with GNN integrated with Transformer or an ensemble of several different models including Transformer, GNN and RNN, a pure Transformer structure is easier for refining, maintaining and testing. And integrating other models with this model would be much easier. In this spirit, we build ISeeCube based on the $2^{\textrm{nd}}$ place solution, by replacing BEiT v2 with TorchScale (the backbone of BEiT-3). We also uses Register Tokens, improving the performance while only slightly increase the computation cost. As a result, the lines of code drop sharply and a lot of new methods could be tested by simply adjusting the config or using other models as the first layer. When having relatively same amount of total parameters, our model outperforms the $2^{\textrm{nd}}$ place solution.
@@
We use rather primitive loss functions such as MSE loss and CE loss in order to explore the fundamental models for Neutrino Events Reconstruction. We then investigated how azimuthal and zenithal distribution would be effected by the choice of regression model or classification model. We show that Transformer, especially Vision Transformer has the potential to learn the information of nodes and time series in neutrino events. Future upgrades on Transformer Encoder or Vision Transformer can be used for Neutrino Events Reconstruction.

### v4

@@ 3.3 Multilayer Perceptron: Regression and Classification
For a continuous space, regression and classification are commonly used methods to predict the output of a neural network. We use different Multilayer Perceptrons (MLPs) as the final layer to build these two models:
@@
  • Regression: The model predicts the output itself. The output is produced by a MLP that results with a tensor with shape $[B,3]$, where $B$ is mini-batch size and 3 represents $x_{pred}$, $y_{pred}$, $z_{pred}$.
@@
We trained three models (The hyper-parameter is summarized in table 4. For detailed training configuration and loss curve, please see section 4.2.):
@@
  • Model S-RegB is a regression model using similar structure, replacing one encoder with two encoders with and without Relative Position Bias, and residual connections are used between these two encoders.
@@
3.4 Loss
We trained the regression model using Mean Squared Error (MSE) Loss, and the classification model using Cross Entropy (CE) Loss. The Competition Metric is used as metric in the validation process and as loss function in the final training epochs.
For regression models, Mean Squared Error (MSE) Loss calculates mean squared error between each element in the prediction $x$ and true value $y$:
$\textrm{MSE}\left(x,y\right)=\frac{1}{n}\sum_{i}\left(x_{i}-y_{i}\right)^{2}$ (3.2)
where $x_{i}$ and $y_{i}$ can be in any shape. In our case, $x_{i}$ and $y_{i}$ are both one-dimensional array with length $3$ (representing $x_{\text{reco}}$, $y_{\text{reco}}$, $z_{\text{reco}}$ and $x_{\text{true}}$, $y_{\text{true}}$, $z_{\text{true}}$ respectively). Also, to prevent exploding gradients, it could be replaced with Smooth L1 Loss [30].
As for classification models, following the Kaggle solutions stated in section 2, we use another commonly used loss function Cross Entropy (CE) Loss:
$\textrm{CE}\left(x,y\right)=\frac{1}{n}\sum_{i}x_{i}\log{y_{i}}$ (3.3)
@@ 3.4 Loss
Mean Angular Error, used as the Kaggle competition metric [10], is the mean arccosine of the inner product of the true angle and the predicted angle:
$\Psi=\arccos(\sin{\vartheta_{\text{true}}}\sin{\vartheta_{\text{reco}}}(\cos{{\varphi_{\text{ture}}}}\cos{{\varphi_{\text{reco}}}}+\sin{{\varphi_{\text{ture}}}}\sin{{\varphi_{\text{reco}}}})+\cos{\vartheta_{\text{true}}}\cos{\vartheta_{\text{reco}}})$ (3.4)
Random guessing will yield a result of $\frac{\pi}{2}$, when the two vectors are orthogonal.
4 Experiments
In section 4.1 we summarize some of the most important aspects of the dataset and the normalization of neutrino events. In section 4.2 we show the learning rate schedule, loss curve and performance of our models, followed by section 4.3 we further explore the performance of our models by comparing azimuthal and zenithal distribution, and propose a new metric: Overlap Ratio.
4.1 Dataset
@@
Each event in the dataset is composed of a series of pulses. Each pulse has $6$ properties: DOM ID (which corresponds to position $x$, $y$, $z$), time, charge and auxiliary.
@@
The biggest challenge of the dataset is its bigness (more than $138$ million events distributed in $660$ batches with the size of over $100$ GBs), so we use the same chunk-based loading method used in the $2^{\textrm{nd}}$ place solution, meaning that every epoch is trained on $654$ batches divided into $8$ sub-epochs, see section 4.2. The advantage of this chunk-based method is that now we can train the model on the entire dataset with relatively low memory cost. However it may have a disadvantage that a random sampling can only sampling across the sub-set of the entire dataset, which may leads to bias. For inference, we used the same batch-based loading method, meaning that we only load the desired batch. This is more flexible.
Figure 3: This is an example event where auxiliary False and True are shown separately. It is intuitive that False will get better result in terms of shape thus direction reconstruction. This figure is originally used in [10].
4.2 Training and Performance
@@
Table 4: Hyper-parameters and score of ISeeCube. “S-RegA”, “S-RegB” and “S-ClsB” refers to different “Small” models (see section 3.3). The score for competition metric is the smaller the better.
@@
Model S-RegA follows the learning rate schedule as table 5. Model S-RegB and S-ClsB follows the learning rate schedule as table 6. The reason for epoch 0 in table 5 is that we first train our model for 1 epoch to test the performance and we think loading this pre-trained model for further training won’t lead to overfitting, considering that (a) the score now is relatively high and (b) there are following 7 epochs with different dataset (see chunk-based method mentioned in section 4.1). In order to save time and resources when training, for epoch 0$\sim$48, we use $N=196$ when loading the data (the pulses in an event is 196 after normalization), and for epoch 49$\sim$56, we use $N=256$. In table 5 and table 6, “COMB” loss means combining two losses by adding them up:
$COMB$ $=\Psi+0.05\cdot MSE$ (4.1)
@@
where $\Psi$ is competition metric in the Kaggle competition. In table 5 and table 6, “div” and “div_final” are hyper-parameter used in Leaner.fit_one_cycle with 1cycle police [31] in fastai.
@@
The loss curve of S-RegA, S-RegB and S-ClsB are shown in figure 4. As we can see, the training loss curve is flucuating while the validation loss curve is constantly reducing. Also, by replacing MSE loss or CE loss with COMB loss, the loss curve drops in a single epoch.
As shown in table 3 and table 4, by adding merely 2.4M parameters, our best model S-RegA outperforms $2^{\textrm{nd}}$ place solution’s S d32 model and approaches its B d32 model. Also, compared with CV512, our model uses CV256, which means that our model requires about half the GPU memory when inferencing.
@@
Figure 5: The distribution of competition metric score of batch 655. By using voting rule as the ensemble method, our model reaches a score of 0.9645.
4.3 Azimuthal and Zenithal Distribution
$\rho=\frac{\text{overlap area of the prediction and the true value}}{\text{area of the true value}}$ (4.2)
As shown in figure 6 (a), for azimuth prediction, the overlap ratio of both regression model and classification model are over 90%. It’s possible to use overlap ratio as loss function, with bigger batch size when training.
@@ 4.3 Azimuthal and Zenithal Distribution
On the contrary, the zenithal distribution is irregular and has lower overlap ratio. The predictions made by regression models tend to be drifted towards $\theta=0$, while the predictions made by classification models tend to be drifted towards $\theta=\frac{\pi}{3}$ and $\theta=\frac{\pi}{2}$. Regression models have a higher overlap ratio than classification models.
@@
As we can see from table 4 and figure 6, this ratio uniformly reflects the performance of the model. By fitting the score of the competition metric and the overlap ratio with four data points (S-RegA, S-RegB, S-ClsB and ensemble), we can get figure 7. The curve for azimuth ($\varphi$) is more linear than the curve for zenith ($\theta$), this is most likely due to the fact that azimuth distribution is more uniform than zenith distribution, as shown in figure 6.
@@
We use some of the most common loss in machine learning to investigate the fundamental models for Neutrino Events Reconstruction. Regression model and Classification model both have their own bias on certain ranges and we are still looking for a unified model. By using ensemble, we can lower the competition metric score, but the distribution of zenith is largely close to regression. Thus ensemble is not the unified model.
@@
Figure 7: Fitting competition metric score with overlap ratio. The observed data indicates that there is a positive correlation between the overlap ratio and the improvement of competition metric score.
5 Discussion and Conclusion
@@
Neutrino Events Reconstruction has always been crucial for neutrino observatories such as IceCube and KM3NeT/ORCA, which traditionally uses CNNs or GNNs. In the Kaggle competition, many solutions use Transformers. And it is shown that a pure Transformer model performs really well for Neutrino Events Reconstruction.
@@
Compared with GNN integrated with Transformer or an ensemble of several different models including Transformer, GNN and RNN, a pure Transformer structure is easier for refining, maintaining and testing. And integrating other models with this model would be much easier. In this spirit, we build ISeeCube based on the $2^{\textrm{nd}}$ place solution, by replacing BEiT v2 with TorchScale (the backbone of BEiT-3). We also uses Register Tokens, improving the performance while only slightly increase the computation cost. The lines of code drop sharply and a lot of new methods could be tested by simply adjusting the config or using other models as the first layer. When having relatively same amount of total parameters, our model outperforms the $2^{\textrm{nd}}$ place solution.
@@
We use rather primitive loss functions such as MSE loss and CE loss in order to explore the fundamental models for Neutrino Events Reconstruction. We then investigated how azimuthal and zenithal distribution would be effected by the choice of regression model or classification model. We show that Transformer, especially Vision Transformer has the potential to learn the information of graphs and time series in neutrino events. Future upgrades on Transformer Encoder or Vision Transformer can be used for Neutrino Events Reconstruction.